# UniPruning: Unifying Local Metric and Global Feedback for Scalable Sparse LLMs

### v1

@@ 1 Introduction
Large Language Models (LLMs) (Achiam et al., 2023; Touvron et al., 2023; Zhang et al., 2022) have redefined the frontier of natural language processing, achieving unprecedented capabilities across diverse tasks. Yet, their deployment at scale remains constrained by prohibitive computational and memory costs driven by their enormous parameter counts. To bridge this gap, model compression has emerged as a critical direction, with quantization (Lin et al., 2024), distillation (Gou et al., 2021), and pruning (Frantar & Alistarh, 2023) as key strategies. Among these, pruning stands out for its ability to induce sparsity while preserving architectural flexibility, thereby delivering substantial reductions in both memory footprint and computational demand.
@@
Existing pruning paradigms for LLMs differ along two axes: structural granularity and algorithmic coordination. Structurally, pruning ranges from unstructured pruning, which removes individual weights for fine-grained control but suffers from limited hardware acceleration, to structured pruning, which eliminates entire channels or neurons to enable efficient execution on modern accelerators. Semi-structured pruning (Mishra et al., 2021), such as the widely adopted N:MN{:}M format, strikes a practical balance, enabling substantial sparsity with hardware-friendly patterns.
@@
From an algorithmic perspective, pruning methods fall into two categories: local metric and global feedback. Local approaches, such as SparseGPT (Frantar & Alistarh, 2023) and Wanda (Sun et al., 2024), make layer-wise pruning decisions based on weight and activation statistics, offering simplicity but often failing under high sparsity due to ignored cross-layer dependencies. Global feedback methods address this by introducing model-wide coordination through regularization or mask learning, as seen in SparseLLM (Bai et al., 2024) and ProxSparse (Liu et al., 2025). While more consistent, these approaches can be computationally costly or restricted by specific sparsity formats.
@@
In this paper, we introduce UniPruning, a unified pruning framework that combines the speed of local metrics with the consistency of global feedback, all without requiring weight updates. To integrate these two objectives, we adopt the mirror descent algorithm as a principled approach for joint optimization. UniPruning employs a fast, layer-wise scoring step to extract local evidence, coupled with a lightweight global controller that redistributes a single sparsity budget across layers using a mirror-descent projection (Beck & Teboulle, 2003; Nemirovsky & Yudin, 1983). Concretely, model weights evolve along a gradient flow while an auxiliary saliency variable $\Gamma$ is updated under a sparsity-aware projection. This mechanism naturally supports both unstructured and semi-structured (N:MN{:}M) pruning within one framework. After calibration, masks are generated by a single sorting operation on $\Gamma$ and directly applied to the original pretrained weights, enabling one-shot extraction of multiple sparsity levels. To stabilize pruning decisions, UniPruning incorporates local saliency signals from a calibration set as robust local signals, while the global controller ensures balanced allocation across layers. This synergy yields pruning that is as efficient as local methods, yet globally consistent and structurally aware like feedback-based approaches.
@@ 2 Related Work
LLM pruning. The tension between local efficiency and global coordination becomes even more pronounced when scaling pruning to large language models. On the one hand, lightweight local approaches such as Wanda (Sun et al., 2024), RIA (Zhang et al., 2024), stochRIA (Yi & Richtárik, 2025) show that activation-aware (relative activation-aware) scoring can prune both unstructured and semi-structured (N:MN{:}M) patterns effectively, requiring only a small calibration set and minimal computation. On the other hand, methods like SparseGPT (Frantar & Alistarh, 2023) push pruning into the LLM regime by incorporating approximate second-order information: they prune many weights in one shot and apply a local least-squares correction to stabilize outputs. These advances confirm that post-training pruning can achieve strong efficiency–accuracy trade-offs at LLM scale, even without retraining. However, they remain fundamentally layer-local: each layer is pruned largely in isolation, with limited capacity to balance sparsity across the entire model (Ding et al., 2025b). This gap highlights the central open question for LLM pruning: how to unify the speed and practicality of local methods with the robustness and balance of global coordination, especially under extreme sparsity or hardware-specific formats.
Mirror descent. Mirror descent (Nemirovsky & Yudin, 1983; Beck & Teboulle, 2003; Bubeck, 2015; Ding et al., 2025a) is a general framework for constrained, geometry-aware optimization. It maps parameters into a dual space, takes gradient steps there, and projects back via a mirror map (Bregman projection). With a Euclidean map it reduces to projected gradient descent; with nonsmooth regularizers it yields proximal updates. This is particularly useful for pruning, where sparsity can be expressed as a regularizer (e.g., $\ell_{1}$, group, or block constraints) and enforced through a proximal step—stabilizing optimization while respecting structural constraints. Mirror descent also connects to continuous-time dynamics in differential-inclusion-based pruning, viewing proximal updates as discrete flows toward sparse sets. We build on this toolkit to unify local saliency with global sparsity budgets in an efficient, structure-aware pruning framework.
@@ 3 Problem Setup
A calibration set $\mathcal{C}$ is typically used to guide the choice of $M$, ensuring that the pruned model behaves similarly to the dense model at the desired sparsity and structure. Such calibration sets are often drawn from common pretraining corpora such as C4 (Raffel et al., 2020), WikiText (Merity et al., 2016), or PTB (Marcinkiewicz, 1994).
@@
Let $\mathcal{D}$ measure the discrepancy between the pruned and dense models on $\mathcal{C}$, and let $\mathcal{C}(M)$ be a global cost (e.g., nonzeros, FLOPs). A budgeted global objective is
$\min_{M\in\mathcal{M}}\ \frac{1}{|\mathcal{C}|}\sum_{x\in\mathcal{C}}\mathcal{D}\!\big(f(x;W_{0}\odot M),\ f(x;W_{0})\big)\quad\text{s.t.}\quad\mathcal{C}(M)\leq B,$ (1)
@@
$\min_{M\in\mathcal{M}}\ \frac{1}{|\mathcal{C}|}\sum_{x\in\mathcal{C}}\mathcal{D}\!\big(f(x;W_{0}\odot M),\ f(x;W_{0})\big)\;+\;\lambda\,\mathcal{C}(M).$ (2)
@@
Local metric pruning. Local metric pruning relies on simple heuristics, such as weight magnitude or weight activation products, to prune parameters independently within each layer. This makes it efficient and easy to apply without heavy global optimization or modifications to pretrained weights. It naturally supports multiple sparsity patterns (unstructured and semi-structured) and can operate in a single pass with only a small calibration set. However, this layer-wise independence comes at a cost: it ignores cross-layer dependencies and trade-offs, which can lead to suboptimal sparse structures and degraded model-level accuracy, particularly at high sparsity. Moreover, many criteria remain heuristic and lack a unified optimization framework, limiting principled control over sparsity allocation. So, its each layer $\ell$ selects a mask under its own budget $B_{\ell}$:
$\min_{M_{\ell}\in\mathcal{M}_{\ell}}\ \frac{1}{|\mathcal{C}|}\sum_{x\in\mathcal{C}}\mathcal{D}_{\ell}\!\big(g_{\ell}(x;W_{0,\ell}\odot M_{\ell}),\ g_{\ell}(x;W_{0,\ell})\big)\quad\text{s.t.}\quad\mathcal{C}_{\ell}(M_{\ell})\leq B_{\ell}.$ (3)
@@ 4 UniPruning
6:  $g_{\text{task}}\leftarrow\nabla_{W}\mathcal{L}_{\text{task}}(W^{n})$
7:  $g_{\text{align}}\leftarrow\rho\cdot\nabla_{W}\tfrac{1}{2}\|\Gamma^{n}-S(W^{n},X)\|_{F}^{2}$
8:  $W^{n+1}\leftarrow W^{n}-\kappa\alpha_{n}\cdot(g_{\text{task}}+g_{\text{align}})$ $\triangleright$ Gradient step on $W$
9:  if N:MN\!:\!M pruning then
10:   $W^{n+1}\leftarrow\operatorname{Prox}_{R_{2:4}}(W^{n+1})$
11:   $R_{2:4}(w)=|w_{1}||w_{2}||w_{3}|+|w_{2}||w_{3}||w_{4}|+|w_{3}||w_{4}||w_{1}|+|w_{4}||w_{1}||w_{2}|$
12:  $V^{n+1}\leftarrow V^{n}-\alpha_{n}\rho\cdot(\Gamma^{n}-S(W^{n},X))$
13:  $\Gamma^{n+1}\leftarrow\operatorname{Prox}_{\Omega}(V^{n+1})$ $\triangleright$ Proximal update on $\Gamma$
@@
15:Export (unstructured): Sort $|\Gamma^{\star}|$ once. For any global budget $B$, set threshold $\tau(B)$ to keep the top-$B$ entries and define mask $\widehat{M}(B)=\mathbb{I}(|\Gamma^{\star}|\geq\tau(B))$. Return $\widetilde{W}(B)=W_{0}\odot\widehat{M}(B)$.
16:Export (N:MN\!:\!M): In each block of size $M$, keep the top-$N$ entries by $|\Gamma^{\star}|$ and zero out the rest. Return $\widetilde{W}_{N\!:\!M}=W_{0}\odot\widehat{M}_{N\!:\!M}$.
@@
This design follows the local metric approach used in Wanda (Sun et al., 2024), and can also incorporate other local methods such as RIA (Zhang et al., 2024); additional experimental comparisons are provided in the appendix.
@@
where $\Omega$ is a sparsity-inducing term. The second term injects local metric by aligning $\Gamma$ with $S(W,X)$; the third term imposes a global sparsity objective via $\Omega$, guiding the gradual gradient updates to $\Gamma$ and $W$ with global feedback. Because $\Omega$ can be non-differentiable such as $L_{1}$, we do not minimize equation 4 directly. Instead, we use a mirror-descent splitting that leads to the following dynamics.
Dynamics and updates. We now present the detailed algorithmic formulation of our proposed pruning method, UniPruning. During the sparsity training stage, we update the model weights $W$ and the saliency scores $\Gamma$ through a coupled dynamic process. The weights $W$ follow a smooth gradient descent, while $\Gamma$ is updated via a proximal step guided by both local activation statistics and a global sparsity constraint.
@@ 4.3 Convergence
Experimental Setup. We evaluate on representative LLM families, including LLaMA2 (Touvron et al., 2023), Qwen2.5 (Team, 2025), and Llama3 (Grattafiori et al., 2024) series, as well as distilled DeepSeek model (Guo et al., 2025). We consider several representative post-training pruning competitors which involve no weight update during pruning, keeping the same with our method: (1) Magnitude (Zhu & Gupta, 2017), the most prevalent pruning approach; (2) Wanda (Sun et al., 2024), which ranks weights using local metric scores and is applicable to both unstructured and semi-structured settings; (3) RIA (Zhang et al., 2024), which combines relative importance with activation norms to provide stable pruning decisions across different sparsity levels; and (4) ProxSparse (Liu et al., 2025), a proximal optimization framework that specifically targets semi-structured N:MN{:}M pruning and achieves state-of-the-art results under 2:4 pattern.
For fairness, we adopt common calibration setup of 128 randomly sampled C4 datasets (Raffel et al., 2020). Model quality is evaluated on both zero-shot reasoning tasks and language modeling: zero-shot performance is measured with the EleutherAI LM-Eval-Harnesss (Gao et al., 2024) on standard benchmark, while WikiText perplexity is reported as the language modeling metric (Merity et al., 2016). For unstructured pruning we use stochRIA (Yi & Richtárik, 2025) as the local metric, and for 2:4 semi-structured pruning we use Wanda (Sun et al., 2024). In both settings, we apply an additional $\lambda L_{1}$ regularization term $\Omega$ with $\lambda=0.001$ (further discussed in Appendix A.3.3). Ablations of the local metric are provided in Section 5.3. For all LLMs, we fix the context length to 4096. All experiments are conducted on a single NVIDIA H200 GPU with 141GB of memory, using a learning rate of 1e-4.
@@ 5.1 Unstructured Pruning
We evaluate unstructured pruning at 60% sparsity on six pretrained LLMs, comparing UniPruning against Magnitude, Wanda, and RIA under the standard 128-sample C4 calibration. We report WikiText perplexity and zero-shot accuracy on ARC-C/E, HellaSwag, OBQA, PIQA, and SIQA (Gao et al., 2024), and further include the average accuracy over all the evaluated benchmarks. Detailed results shows in Table 1.
@@ 5.2 Semi-Structured N:MN\!:\!M Pruning
For semi-structured pruning, we primarily evaluate the 2:4 sparsity pattern across models. To align with this hardware-friendly pattern, we incorporate a 2:4 regularizer into our algorithm. The detailed formulation and implementation of this adaptation are provided in Algorithm 1.
@@
Beyond perplexity, we also evaluate zero-shot performance on downstream benchmarks under the 2:4 constraint, with results provided in Appendix A.2. Furthermore, to situate our approach against stronger baselines, we conduct an additional comparison with SparseGPT (Frantar & Alistarh, 2023), which allows weight updates during pruning. The corresponding results are also reported in Appendix A.2.

### v2

@@ 1 Introduction
Large Language Models (LLMs) (achiam2023gpt; touvron2023llama; zhang2022opt) have redefined the frontier of natural language processing, achieving unprecedented capabilities across diverse tasks. Yet, their deployment at scale remains constrained by prohibitive computational and memory costs driven by their enormous parameter counts. To bridge this gap, model compression has emerged as a critical direction, with quantization (lin2024awq), distillation (gou2021knowledge), and pruning (frantar2023sparsegpt) as key strategies. Among these, pruning stands out for its ability to induce sparsity while preserving architectural flexibility, thereby delivering substantial reductions in both memory footprint and computational demand.
@@
Existing pruning paradigms for LLMs differ along two axes: structural granularity and algorithmic coordination. Structurally, pruning ranges from unstructured pruning, which removes individual weights for fine-grained control but suffers from limited hardware acceleration, to structured pruning, which eliminates entire channels or neurons to enable efficient execution on modern accelerators. Semi-structured pruning (mishra2021accelerating), such as the widely adopted N:MN{:}M format, strikes a practical balance, enabling substantial sparsity with hardware-friendly patterns.
@@
From an algorithmic perspective, pruning methods fall into two categories: local metric and global feedback. Local approaches, such as SparseGPT (frantar2023sparsegpt) and Wanda (sun2024simple), make layer-wise pruning decisions based on weight and activation statistics, offering simplicity but often failing under high sparsity due to ignored cross-layer dependencies. Global feedback methods address this by introducing model-wide coordination through regularization or mask learning, as seen in SparseLLM (bai2024sparsellm) and ProxSparse (liu2025proxsparse). While more consistent, these approaches can be computationally costly or restricted by specific sparsity formats.
@@
In this paper, we introduce UniPruning, a unified pruning framework that combines the speed of local metrics with the consistency of global feedback, all without requiring weight updates. To integrate these two objectives, we adopt the mirror descent algorithm as a principled approach for joint optimization. UniPruning employs a fast, layer-wise scoring step to extract local evidence, coupled with a lightweight global controller that redistributes a single sparsity budget across layers using a mirror-descent projection (beck2003mirror; nemirovsky1983problem). Concretely, model weights evolve along a gradient flow while an auxiliary saliency variable $\Gamma$ is updated under a sparsity-aware projection. This mechanism naturally supports both unstructured and semi-structured (N:MN{:}M) pruning within one framework. After calibration, masks are generated by a single sorting operation on $\Gamma$ and directly applied to the original pretrained weights, enabling one-shot extraction of multiple sparsity levels. To stabilize pruning decisions, UniPruning incorporates local saliency signals from a calibration set as robust local signals, while the global controller ensures balanced allocation across layers. This synergy yields pruning that is as efficient as local methods, yet globally consistent and structurally aware like feedback-based approaches.
@@ 2 Related Work
LLM pruning. The tension between local efficiency and global coordination becomes even more pronounced when scaling pruning to large language models. On the one hand, lightweight local approaches such as Wanda (sun2024simple), RIA (zhang2024plug), stochRIA (yi2025symmetric) show that activation-aware (relative activation-aware) scoring can prune both unstructured and semi-structured (N:MN{:}M) patterns effectively, requiring only a small calibration set and minimal computation. On the other hand, methods like SparseGPT (frantar2023sparsegpt) push pruning into the LLM regime by incorporating approximate second-order information: they prune many weights in one shot and apply a local least-squares correction to stabilize outputs. These advances confirm that post-training pruning can achieve strong efficiency–accuracy trade-offs at LLM scale, even without retraining. However, they remain fundamentally layer-local: each layer is pruned largely in isolation, with limited capacity to balance sparsity across the entire model (ding2025revisiting). This gap highlights the central open question for LLM pruning: how to unify the speed and practicality of local methods with the robustness and balance of global coordination, especially under extreme sparsity or hardware-specific formats.
Mirror descent. Mirror descent (nemirovsky1983problem; beck2003mirror; bubeck2015convex; ding2025adaptive) is a general framework for constrained, geometry-aware optimization. It maps parameters into a dual space, takes gradient steps there, and projects back via a mirror map (Bregman projection). With a Euclidean map it reduces to projected gradient descent; with nonsmooth regularizers it yields proximal updates. This is particularly useful for pruning, where sparsity can be expressed as a regularizer (e.g., $\ell_{1}$, group, or block constraints) and enforced through a proximal step—stabilizing optimization while respecting structural constraints. Mirror descent also connects to continuous-time dynamics in differential-inclusion-based pruning, viewing proximal updates as discrete flows toward sparse sets. We build on this toolkit to unify local saliency with global sparsity budgets in an efficient, structure-aware pruning framework.
@@ 3 Problem Setup
A calibration set $\mathcal{C}$ is typically used to guide the choice of $M$, ensuring that the pruned model behaves similarly to the dense model at the desired sparsity and structure. Such calibration sets are often drawn from common pretraining corpora such as C4 (raffel2020exploring), WikiText (merity2016pointer), or PTB (marcinkiewicz1994building).
@@
Let $\mathcal{D}$ measure the discrepancy between the pruned and dense models on $\mathcal{C}$, and let $Cost(M)$ be a global cost (e.g., nonzeros, FLOPs). A budgeted global objective is
$\min_{M\in\mathcal{M}}\ \frac{1}{|\mathcal{C}|}\sum_{x\in\mathcal{C}}\mathcal{D}\!\big(f(x;W_{0}\odot M),\ f(x;W_{0})\big)\quad\text{s.t.}\quad Cost(M)\leq B,$ (1)
@@
$\min_{M\in\mathcal{M}}\ \frac{1}{|\mathcal{C}|}\sum_{x\in\mathcal{C}}\mathcal{D}\!\big(f(x;W_{0}\odot M),\ f(x;W_{0})\big)\;+\;\lambda\,Cost(M).$ (2)
@@
Local metric pruning. Local metric pruning relies on simple heuristics, such as weight magnitude or weight activation products, to prune parameters independently within each layer. This makes it efficient and easy to apply without heavy global optimization or modifications to pretrained weights. It naturally supports multiple sparsity patterns (unstructured and semi-structured) and can operate in a single pass with only a small calibration set. However, this layer-wise independence comes at a cost: it ignores cross-layer dependencies and trade-offs, which can lead to suboptimal sparse structures and degraded model-level accuracy, particularly at high sparsity. Moreover, many criteria remain heuristic and lack a unified optimization framework, limiting principled control over sparsity allocation. So, its each layer $\ell$ selects a mask under its own budget $B_{\ell}$, where $g_{\ell}$ represents the output of layer $\ell$ :
$\min_{M_{\ell}\in\mathcal{M}_{\ell}}\ \frac{1}{|\mathcal{C}|}\sum_{x\in\mathcal{C}}\mathcal{D}_{\ell}\!\big(g_{\ell}(x;W_{0,\ell}\odot M_{\ell}),\ g_{\ell}(x;W_{0,\ell})\big)\quad\text{s.t.}\quad Cost_{\ell}(M_{\ell})\leq B_{\ell}.$ (3)
@@ 4 UniPruning
6:  $S(W^{n},X)\leftarrow\text{Local metric at current }W^{n}$ $\triangleright$ Recompute local statistics every iteration
7:  $g_{\text{task}}\leftarrow\nabla_{W}\mathcal{L}_{\text{task}}(W^{n})$
8:  $g_{\text{align}}\leftarrow\rho\cdot\nabla_{W}\tfrac{1}{2}\|\Gamma^{n}-S(W^{n},X)\|_{F}^{2}$
9:  $W^{n+1}\leftarrow W^{n}-\kappa\alpha_{n}\cdot(g_{\text{task}}+g_{\text{align}})$ $\triangleright$ Gradient step on $W$
10:  if N:MN\!:\!M pruning then
11:   $W^{n+1}\leftarrow\operatorname{Prox}_{R_{2:4}}(W^{n+1})$
12:   $R_{2:4}(w)=|w_{1}||w_{2}||w_{3}|+|w_{2}||w_{3}||w_{4}|+|w_{3}||w_{4}||w_{1}|+|w_{4}||w_{1}||w_{2}|$
13:  $V^{n+1}\leftarrow V^{n}-\alpha_{n}\rho\cdot(\Gamma^{n}-S(W^{n},X))$
14:  $\Gamma^{n+1}\leftarrow\operatorname{Prox}_{\Omega}(V^{n+1})$ $\triangleright$ Proximal update on $\Gamma$
@@
16:Export (unstructured): Sort $|\Gamma^{\star}|$ once. For any global budget $B$, set threshold $\tau(B)$ to keep the top-$B$ entries and define mask $\widehat{M}(B)=\mathbb{I}(|\Gamma^{\star}|\geq\tau(B))$. Return $\widetilde{W}(B)=W_{0}\odot\widehat{M}(B)$.
17:Export (N:MN\!:\!M): In each block of size $M$, keep the top-$N$ entries by $|\Gamma^{\star}|$ and zero out the rest. Return $\widetilde{W}_{N\!:\!M}=W_{0}\odot\widehat{M}_{N\!:\!M}$.
@@
This design follows the local metric approach used in Wanda (sun2024simple), and can also incorporate other local methods such as RIA (zhang2024plug); additional experimental comparisons are provided in the appendix.
@@
where $\Omega$ is a sparsity-inducing term. The second term injects local metric by aligning $\Gamma$ with $S(W,X)$, $\rho$ as a hyperparameter; the third term imposes a global sparsity objective via $\Omega$, guiding the gradual gradient updates to $\Gamma$ and $W$ with global feedback. Because $\Omega$ can be non-differentiable such as $L_{1}$, we do not minimize equation 4 directly. Instead, we use a mirror-descent splitting that leads to the following dynamics.
Dynamics and updates. We now present the detailed algorithmic formulation of our proposed pruning method, UniPruning. During the sparsity training stage, we update the model weights $W$ and the saliency scores $\Gamma$ through a coupled dynamic process, with $\mathbf{V}$ acts as a conjugate variable (or dual variable). The weights $W$ follow a smooth gradient descent, while $\Gamma$ is updated via a proximal step guided by both local activation statistics and a global sparsity constraint.
@@ 4.3 Convergence
Experimental Setup. We evaluate on representative LLM families, including LLaMA2 (touvron2023llama), Qwen2.5 (qwen2.5), and Llama3 (grattafiori2024llama3) series, as well as distilled DeepSeek model (guo2025deepseek). We consider several representative post-training pruning competitors which involve no weight update during pruning, keeping the same with our method: (1) Magnitude (zhu2017prune), the most prevalent pruning approach; (2) Wanda (sun2024simple), which ranks weights using local metric scores and is applicable to both unstructured and semi-structured settings; (3) RIA (zhang2024plug), which combines relative importance with activation norms to provide stable pruning decisions across different sparsity levels; and (4) ProxSparse (liu2025proxsparse), a proximal optimization framework that specifically targets semi-structured N:MN{:}M pruning and achieves state-of-the-art results under 2:4 pattern.
For fairness, we adopt common calibration setup of 128 randomly sampled C4 datasets (raffel2020exploring). Model quality is evaluated on both zero-shot reasoning tasks and language modeling: zero-shot performance is measured with the EleutherAI LM-Eval-Harnesss (eval-harness) on standard benchmark, while WikiText perplexity is reported as the language modeling metric (merity2016pointer). For unstructured pruning we use stochRIA (yi2025symmetric) as the local metric, and for 2:4 semi-structured pruning we use Wanda (sun2024simple). In both settings, we apply an additional $\lambda L_{1}$ regularization term $\Omega$ with $\lambda=0.001$ (further discussed in Appendix A.3.3). Ablations of the local metric are provided in Section 5.4. For all LLMs, we fix the context length to 4096. All experiments are conducted on a single NVIDIA H200 GPU with 141GB of memory, using a learning rate of 1e-4.
@@ 5.1 Unstructured Pruning
We evaluate unstructured pruning at 60% sparsity on six pretrained LLMs, comparing UniPruning against Magnitude, Wanda, and RIA under the standard 128-sample C4 calibration. We report WikiText perplexity and zero-shot accuracy on ARC-C/E, HellaSwag, OBQA, PIQA, and SIQA (eval-harness), and further include the average accuracy over all the evaluated benchmarks. Detailed results shows in Table 1.
@@ 5.2 Semi-Structured N:MN\!:\!M Pruning
For semi-structured pruning, we primarily evaluate the 2:4 sparsity pattern across models. To align with this hardware-friendly pattern, we incorporate a 2:4 regularizer  (kübler2025proximaloperatorinducing24sparsity) into our algorithm. The detailed formulation and implementation of this adaptation are provided in Algorithm 1.
@@
Beyond perplexity, we also evaluate zero-shot performance on downstream benchmarks under the 2:4 constraint, with results provided in Appendix A.2. Furthermore, to situate our approach against stronger baselines, we conduct an additional comparison with SparseGPT (frantar2023sparsegpt), which allows weight updates during pruning. The corresponding results are also reported in Appendix A.2.
@@
5.3 Results On Ascend NPU
We additionally evaluate openPangu architecture on Huawei Ascend NPU, as reported in Table 3. This model adopts a substantially different architecture and training paradigm from other models, making it a compelling cross-architecture validation case. Under the unstructured 50% sparsity and 2:4 sparsity setting, UniPruningall significantly outperforms baselines on openPangu-7B and interestingly finds the winning structure of the lottery ticket. This demonstrates that UniPruning generalizes robustly to different model architectures and also heterogeneous hardware backends (GPU/NPU), providing concrete empirical evidence for its architectural agnosticism.
Table 3: Results On Ascend NPU.
[…]
openPangu-Embedded-7B-V1.1 dense 31.36 0.3302 0.5673 0.3946 0.4497 0.1980 0.6844 0.4374
2:4 ria 208.04 0.2526 0.5034 0.3792 0.3645 0.1600 0.6344 0.3824
2:4 wanda 237.32 0.2628 0.5109 0.3746 0.3593 0.1600 0.6328 0.3834
2:4 UniPruning 106.21 0.2927 0.6002 0.3930 0.3778 0.1840 0.6518 0.4166
50% ria 59.75 0.3072 0.5488 0.3936 0.4211 0.1960 0.6632 0.4217
50% wanda 70.26 0.2944 0.5492 0.3930 0.4227 0.2060 0.6670 0.4221
50% UniPruning 49.73 0.3677 0.7054 0.4043 0.6700 0.2040 0.6959 0.5079